OpenStudy (anonymous):

find a polar representation for the curve: x^2 + y^2 = 9

myininaya (myininaya):

do you remember that \[r^2=x^2+y^2\]

OpenStudy (anonymous):

yeah so r = 3

myininaya (myininaya):

satellite I having a moment here doesn't r=3 include the equation r=-3?

myininaya (myininaya):

you know when we are talking about polar equations?

myininaya (myininaya):

i would include r=-3 just in case because i'm having a memory issue right now

OpenStudy (anonymous):

i think it is just \[r=9\]

myininaya (myininaya):

\[r^2=9 => r=\pm 3\]

OpenStudy (anonymous):

r is the radius, always non negative. you want \[r=f(\theta) but here r is constant

OpenStudy (anonymous):

\[r=f(\theta)\]

myininaya (myininaya):

you only need r=3

OpenStudy (anonymous):

right i lunched it is \[r=3\]

OpenStudy (anonymous):

\[x=rcos \theta\] \[y=rsin \theta\] r=3\[r ^{2}\cos ^{2}\theta+r ^{2}\sin ^{2}\theta=9\]

OpenStudy (anonymous):

yeah i got that far

OpenStudy (anonymous):

yeah but this says \[r=3\]

OpenStudy (anonymous):

how do i simplify that?

OpenStudy (anonymous):

too much work. r is the radius. it is a constant since you have a circle of radius 3

myininaya (myininaya):

factor out r^2

OpenStudy (anonymous):

did that

myininaya (myininaya):

cos^2(theta)+sin^2(theta)=1

OpenStudy (anonymous):

\[\cos ^{2}\theta+\sin ^{2}\theta=1\]

OpenStudy (anonymous):

okay

myininaya (myininaya):

you don't have to do it that way the easiest is just recalling \[r^2=x^2+y^2\]

OpenStudy (anonymous):

you shouldn't think that hard! it is true that \[\cos ^{2}\theta+\sin ^{2}\theta=1\] but that is way too much work

OpenStudy (anonymous):

so it's just r = 3 as my answeR?

myininaya (myininaya):

yes

OpenStudy (anonymous):

yes a circle looks like \[r= number\]

myininaya (myininaya):

r=3 will include all points from the center that have distance 3 from it

OpenStudy (anonymous):

in polar coordinates a circle is just r = a number?

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

r after all stands for "radius" and circle is a figure where the radius is constant

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=r+%3D+3+

OpenStudy (anonymous):

here is one where r is not constant http://www.wolframalpha.com/input/?i=r+%3D+1%2Bsin%28theta%29